Old Country Links, Inc., produces sausages in three production departments—Mixing, Casing and Curing, and Packaging. In the Mixing Department, meats are prepared and ground and then mixed with spices. The spiced meat mixture is then transferred to the Casing and Curing Department, where the mixture is force-fed into casings and then hung and cured in climate-controlled smoking chambers. In the Packaging Department, the cured sausages are sorted, packed, and labeled. The company uses the weighted-average method in its process costing system. Data for September for the Casing and Curing Department follow:

Percent Completed
Units Mixing Materials Conversion
Work in process inventory, September 1 6 100 % 60 % 50 %
Work in process inventory, September 30 6 100 % 20 % 10 %
Mixing Materials Conversion
Work in process inventory, September 1 $ 12,204 $ 54 $ 2,796
Cost added during September $ 174,976 $ 14,485 $ 131,474

Mixing cost represents the costs of the spiced meat mixture transferred in from the Mixing Department. The spiced meat mixture is processed in the Casing and Curing Department in batches; each unit in the above table is a batch and one batch of spiced meat mixture produces a set amount of sausages that are passed on to the Packaging Department. During September, 92 batches (i.e., units) were completed and transferred to the Packaging Department.

Required:

1. Determine the Casing and Curing Department's equivalent units of production for mixing, materials, and conversion for the month of September.

2. Compute the Casing and Curing Department's cost per equivalent unit for mixing, materials, and conversion for the month of September.

3. Compute the Casing and Curing Department's cost of ending work in process inventory for mixing, materials, conversion, and in total for September.

4. Compute the Casing and Curing Department's cost of units transferred out to the Packaging Department for mixing, materials, conversion, and in total for September.

5. Prepare a cost reconciliation report for the Casing and Curing Department for September.
Answer #1

1) Equivalent unit

Whole Units % mixing EUP-Mixing % material EUP-Material % Conversion EUP-Conversion
Unit transferred out 92 100% 92 100% 92 100% 92
Ending WIP 6 100% 6 20% 1.2 10% 0.6
Total 98 98 93.2 92.60

2) Cost per equivalent unit

Mixing Material Conversion
Beginning WIP 12204 54 2796
Cost added 174976 14485 131474
Total Cost 187180 14539 134270
Equivalent unit 98 93.2 92.60
Cost per equivalent unit 1910 156 1450

3) Cost of ending WIP

Mixing Material Conversion Total
Cost of ending WIP 6*1910 = 11460 1.2*156 = 187 1450*.6 = 870 12517

4) Cost of units transferred out

Mixing Material Conversion Total
Cost of Unit transferred out 92*1910 = 175720 92*156 = 14352 92*1450 = 133400 323472

5) Cost reconciliation

Cost to be accounted for
Beginning WIP 15054
Cost added 320935
Total cost to be accounted for 335989
Cost accounted as
Cost of unit transferred out 323472
Cost of ending WIP 12517
Total Cost accounted as 335989